OpenStudy (anonymous):

Use cross products to solve the proportion. Must show work. 1. 1/3 = z/15 2. 3/8 = s/16

OpenStudy (math&ing001):

For the first multiply both sides by 15. For the second multiply by 16.

OpenStudy (anonymous):

?

OpenStudy (math&ing001):

A question mark isn't a constructive response. \[\frac{ 1 }{ 3 }=\frac{ z }{ 15 }\] \[\frac{ 1 }{ 3 }*15=\frac{ z }{ 15 }*15\] \[z=\frac{ 15 }{ 3 }=5\] Do the same for the 2nd one.